prove that the tangent to the curve y=(1+3x^2)/(3+x^2) drawn at the point y=1 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
If we put y = 1
 
Then
 
1 = (1+3x²) /(3 +x²)
 
Now
3 + x² = 1 + 3x²
 
2x² = 2
 
x = ± 1
 
Now find dy/dx and then find the equation of tangent